What is second equation of motion.
1
Expert's answer
2020-12-02T07:34:06-0500

The second equation of motion gives the displacement of an object under constant acceleration:

S=V(initial)"\times" t+ "\frac{a \times t^2}{2}" ,

S - travelled distance,

V(initial) - initial velocity,

t - spent time and

a - acceleration.

Let`s to proof it:

We saw above the Second Equation of Motion

S=V(initial)"\times" t+ "\frac{a \times t^2}{2}"

Derivation We know that Velocity = "\frac{Displacement}{Time}"  If Velocity is not constant (i.e. Velocity keeps on increasing or decreasing) We can also take Average Velocity in place of Velocity.

So our formula becomes Displacement = Average Velocity "\times" Time 

Displacement ="\frac{Initial velocity + Final velocity}{2}""\times" Time

S="\frac{V(initial)+V(final)}{2}""\times"t, V(final)- final velocity.

From first equation of motion, we know that: V(final)=V(initial)+a"\times"t, a - acceleration.

Putting value of Vf in this equation:

S="\frac{V(initial)+(V(initial)+a \times t)}{2} \times t",

S=(V(initial)+"\frac{a \times t}{2}")"\times"t

S=V(initial)"\times" t+"\frac{a \times t^2}{2}" . So that is answer!